I have worked this problem as far as I can, but cannot figure out how to solve for the rate of the river.
Question is: If a boat goes downstream 72 miles in 3 hours and upstream 60 miles in 6 hours, the rate of the river and the rate of the boat in still water respectively are ______?
I've figured out that the boat is going 24 mph downstream and 10 mph upstream, but don't know the method to use in calculating the rate of the boat/river in still water.

Answer by macston(5194)

Downstream rate=rate of boat+rate of river
Upstream rate=rate of boat-rate of river
Let b=rate of boat and r=rate of river
24mph=b+r Subtract b from each side
24 mph-b=r
10mph=b-r Substitute for r from above
10 mph=b-(24 mph-b)
10 mph=b-24mph+b Add 24 mph to each side
34 mph=2b Divide each side by 2.
17 mph=b ANSWER 1: The rate of the boat in still water is 17 mph.
r=24mph-b=24mph-17mph=7 mph ANSWER 2:The rate of the river is 7 mph